A manufacturer claims that the average lifetime of his lightbulb is

3 years or 36 months. The standard deviation is 8 months. Fifty bulbs are selected, and the average life expectancy is found to be 32 months. Should the manufacturer statement be rejected at level of significance 0.01?


1
Expert's answer
2022-06-02T08:32:10-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=36"

"H_1:\mu\not=36"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is "z_c = 2.5758."

The rejection region for this two-tailed test is "R = \{z:|z|>2.5758\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{32-36}{8/\sqrt{50}}\approx-3.5355"

Since it is observed that "|z|=3.5355>2.5758=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=2P(z<-3.5355)= 0.000407," and since "p= 0.000407<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 36, at the "\alpha = 0.01" significance level.